# Movable Antennas Meet Intelligent Reflecting Surface: Friends or Foes?

### v1

@@ Abstract
Movable antennas, intelligent reflecting surfaces, near-field communications, performance analysis, alternating optimization, graph theory
@@ I Introduction
On the other hand, movable antenna (MA) (also known as fluid antenna system (FAS) for antenna position adjustment [9]) has drawn great interest in academia and industry recently. Similar to the IRS, MA can reconfigure wireless channel conditions by enabling multiple antennas to be flexibly moved within a confined region at the transmitter or receiver [10, 11, 12]. For example, by optimizing antenna positions, MAs can achieve flexible beamforming for interference nulling and wide-beam coverage [13, 14, 15]. The antenna position optimization for MAs has been investigated in various scenarios, such as physical-layer security [16, 17], multiple-input multiple-output (MIMO) system [18], over-the-air computation [19], cognitive radio [20], integrated sensing and communication (ISAC) [21], etc. Particularly, most of the existing works applied the gradient-based algorithms, e.g., the successive convex approximation (SCA) technique [13, 16, 18, 22] and gradient ascend/descent method [23, 24], to optimize the antenna positions. In addition, some other works proposed a variety of evolutionary algorithms to optimize the antenna positions, such as the particle swarm optimization (PSO) algorithm and its derivative algorithms [19, 25, 26]. Furthermore, the deep learning-based method was introduced in [27] for MA-assisted multicasting, where the feedforward neural networks (FNNs) were utilized to extract useful features from the input direction information and jointly optimize the beamforming and antenna positions. Instead of optimizing the antenna positions in a continuous space, the authors of [28, 20, 17] discretized the movable region into a multitude of sampling points and proposed a graph-based method to select an optimal subset of sampling points for maximizing the communication rates. However, all of the above position optimization algorithms rely on global channel state information (CSI) at any position within the transmit/receive movable region. To facilitate the practical implementation, the authors of [29] proposed a CSI-free MA position optimization approach via the zero-order gradient approximation method.
@@ II System Model and Problem Formulation
Let $\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})\in\mathbb{C}^{M\times N}$ and $\boldsymbol{h}_{\text{IU},k}\in\mathbb{C}^{M\times 1}$ denote the BS-IRS and IRS-user $k$ channels, respectively. Note that the BS-IRS channel depends on the APV, i.e., $\boldsymbol{T}$, while $\boldsymbol{h}_{\text{IU},k}$ is regardless of it. As such, the cascaded BS-IRS-user $k$ channel can be expressed as $\boldsymbol{h}_{k}^{H}(\boldsymbol{T},\boldsymbol{\Phi})=\boldsymbol{h}_{\text{IU},k}^{H}\boldsymbol{\Phi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})\in\mathbb{C}^{1\times N}$, $\forall k\in\mathcal{K}$. Let $\boldsymbol{w}_{k}\in\mathbb{C}^{N\times 1}$ denote the BS’s transmit precoding for user $k$ with $\sum_{k\in\mathcal{K}}{||\boldsymbol{w}_{k}||_{2}}\leq P$, where $P$ is the BS’s transmit power. Then, the received signal at user $k$ is given by
@@
$R_{k}(\boldsymbol{W},\boldsymbol{\Phi},\boldsymbol{T})=\log_{2}\left(1+\frac{\left|\boldsymbol{h}_{k}^{H}(\boldsymbol{\Phi},\boldsymbol{T})\boldsymbol{w}_{k}\right|^{2}}{\sum_{i\neq k}{\left|\boldsymbol{h}_{k}^{H}(\boldsymbol{\Phi},\boldsymbol{T})\boldsymbol{w}_{i}\right|^{2}}+\sigma^{2}}\right)$ (2)
where $\boldsymbol{W}=[\boldsymbol{w}_{1},\boldsymbol{w}_{2},\cdots,\boldsymbol{w}_{K}]\in\mathbb{C}^{N\times K}$ denotes the transmit precoding matrix. Our objective is to maximize the sum-rate of the $K$ users, by jointly optimizing the BS’s transmit precoding matrix $\boldsymbol{W}$, the IRS’s reflection matrix $\boldsymbol{\Phi}$ and the APV $\boldsymbol{T}$. Hence, the optimization problem is formulated as
@@
$||\boldsymbol{t}_{i}-\boldsymbol{t}_{j}||\geq D_{\min},\forall i,j\in\mathcal{N},i\neq j,$ (3c)
@@
where $D_{\min}$ denotes the minimum spacing between any two MAs to avoid mutual coupling. To characterize the performance limit and ease the performance analysis, we assume that all required channel state information is available by applying the existing channel estimation techniques dedicated to MAs/IRSs, e.g., [39] and [40]. In addition, we assume that the antenna movement delay is negligible compared to the channel coherence time, which usually holds in various slow-varying scenarios, e.g., smart homes and factories. We can also reduce the movement delay by applying more advanced electronically-driven solutions for MAs [10].
@@ III-A Performance Analysis
$G_{u}(\boldsymbol{t})=\left(\frac{\lambda}{4\pi}\right)^{2}\left|\sum_{m_{y}=1}^{M_{y}}\sum_{m_{z}=1}^{M_{z}}\frac{|h_{\text{IU},m_{y},m_{z}}|}{D(\boldsymbol{t},m_{y},m_{z})}\right|^{2},$ (9)
@@ III-A1 Optimal Antenna Position
$h_{\text{BU}}(\boldsymbol{t})=\frac{\lambda}{4\pi||\boldsymbol{t}-\boldsymbol{q}_{U}||}e^{j\frac{2\pi}{\lambda}||\boldsymbol{t}-\boldsymbol{q}_{U}||},$ (14)
@@
$\boldsymbol{t}^{\star}=\arg\underset{\boldsymbol{t}\in\mathcal{C}_{t}}{\min}\,\,||\boldsymbol{t}-\boldsymbol{q}_{U}||.$ (15)
@@ III-A2 Fluctuation of Channel Power Gain
On the other hand, we analyze the effects of the BS-IRS distance on (17). For simplicity, we assume that the MA array is perpendicular to the IRS. Then, $R(\boldsymbol{t}_{i})$ can be expressed as $R(\boldsymbol{t}_{i})=\sqrt{d_{\text{BI}}^{2}+d^{2}(\boldsymbol{t}_{i})}$, $i=1,2$, where $d(\boldsymbol{t}_{i})=||\boldsymbol{q}_{B}-\boldsymbol{t}_{i}||$ denotes the distance between the $i$-th location and the center of the MA array, and $d_{\text{BI}}$ denotes the distance from the center of $\mathcal{C}_{t}$ to that of the IRS. As $R(\boldsymbol{t}_{1})<R(\boldsymbol{t}_{2})$, $H(M_{y},M_{z},\boldsymbol{t}_{1})\pm H(M_{y},M_{z},\boldsymbol{t}_{2})$ monotonically decreases with $d_{\text{BI}}$. Thus, $G_{u}(\boldsymbol{t}_{1})-G_{u}(\boldsymbol{t}_{2})$ also decreases with $d_{\text{BI}}$. This indicates that a more/less significant fluctuation of (10) will be resulted if the BS-IRS distance reduces/increases.
@@ III-A3 Far-Field BS-IRS Channel
All of the above analytical results are derived under the general NUSW-based BS-IRS channel model. Next, we consider a special case with the far-field BS-IRS channel, under which the distance between the MA and the $(m_{y},m_{z})$-th IRS reflecting element can be treated as identical, i.e.,
@@
$\boldsymbol{w}(\boldsymbol{T})=\frac{\boldsymbol{v}(\boldsymbol{T})}{||\boldsymbol{v}(\boldsymbol{T})||},$ (19)
@@ III-B Proposed Solution to (P2)
$\boldsymbol{w}(\boldsymbol{T},\boldsymbol{\Phi})=\frac{\left(\boldsymbol{h}_{\text{IU}}^{H}\boldsymbol{\Phi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})\right)^{H}}{||\boldsymbol{h}_{\text{IU}}^{H}\boldsymbol{\Phi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})||}.$ (20)
@@
(P2) $\underset{\boldsymbol{\Phi},\boldsymbol{T}}{\max}\quad\gamma(\boldsymbol{T},\boldsymbol{\Phi})=||\boldsymbol{h}_{\text{IU}}^{H}\boldsymbol{\Phi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})||^{2}$
@@ III-B1 Optimizing $\boldsymbol{\Phi}$ for a Given $\boldsymbol{T}$
(P2-1) $\underset{\boldsymbol{\varphi}}{\max}\quad\gamma(\boldsymbol{T},\boldsymbol{\Phi})=\left|\left|\sum_{m=1}^{M}{\boldsymbol{g}_{1,m}e^{j\varphi_{m}}}\right|\right|^{2}$
@@
which, however, is still a non-convex optimization problem. To tackle this non-convexity, we utilize an element-wise BCD method to optimize the IRS reflecting coefficients sequentially. Specifically, it can be easily shown that for any given $\varphi_{i}$, $i\in{\cal M}$, $i\neq m$, the optimal $\varphi_{m}$ for (P2-1) is given by
$\varphi_{m}^{\star}=\arg\left(\boldsymbol{\alpha}_{m}\boldsymbol{g}_{1,m}^{H}\right),$ (23)
where $\boldsymbol{\alpha}_{m}=\sum_{i\neq m}^{M}{\boldsymbol{g}_{1,i}e^{-j\varphi_{i}}}$. As such, in the $m$-th BCD iteration, we can fix $\varphi_{i}$, $i\neq m$, $i\in{\cal M}$ and optimize $\varphi_{m}$ as (23). Then, the BCD iteration proceeds until the phase shifts of all $M$ IRS reflecting elements have been updated.
@@ IV-B1 Optimizing $\boldsymbol{W}$ for Given $\mathbf{\Phi}$ and $\boldsymbol{T}$
where $\mu\geq 0$ is the optimal dual variable that ensures $\sum_{k=1}^{K}\left\|\boldsymbol{w}_{k}\right\|^{2}\leq P$. Note that $\sum_{k=1}^{K}\left\|\boldsymbol{w}_{k}\right\|^{2}$ is a monotonically decreasing function w.r.t. $\mu$ [33]. Thus, we can find $\mu$ via a bisection search.
@@ V-A Single-User System
First, we consider the single-user setup and plot in Fig. 3 the achievable rate at the user versus the BS-IRS distance, with $N=4$ and $L=8$. The IRS-user distance is fixed as $d_{\text{IU}}=30$ m. It is observed that our proposed algorithm achieves higher received SNRs than all other benchmarks, thanks to channel reconfiguration abilities for both MAs and IRSs. However, the performance gain of our proposed algorithm over the FPA benchmark is marginal and decreases with the BS-IRS distance, which shows a similar trend to the case with the LoS BS-IRS channel in Fig. 2(b). In contrast, if the IRS phase shifts are configured randomly, employing MAs can yield more significant performance gain over the conventional FPAs. For example, when the BS-IRS distance is 3 m, the performance gain of the proposed algorithm over the FPA benchmark is only 0.22 bps/Hz (or 2.01%); while that of the MA-RPS benchmark over the FPA-RPS benchmark is 1.41 bps/Hz (or 39.67%). This is mainly attributed to the channel reconfiguration from the IRS. More specifically, under the optimized IRS reflection, the correlations among all transmit paths between the BS and the IRS may be decreased, which also reduces the spatial diversity gain reaped from the multi-path effect. While under the random IRS reflection, the IRS can be simply treated as a scatterer in the environment, thus barely compromising the multi-path effect.
To validate the above claim, we plot in Figs. 4(a) and 4(b) the channel power gains from the BS to the user within $\mathcal{C}_{t}$ by the proposed algorithm and the MA-RPS benchmark, respectively. It is observed that compared to the proposed algorithm, the MA-RPS benchmark yields more local maxima and minima, i.e., more significant fluctuation, within the transmit array. Furthermore, the gap between the maximum and the minimum channel power gains in the MA-RPS benchmark is around 15.62 dB, which far exceeds that in the proposed algorithm, i.e., 3.48 dB. This observation validates our previous claim regarding the effects of the random versus optimized IRS reflections.

### v2

@@ Abstract
Movable antennas, intelligent reflecting surfaces, near-field communications, performance analysis, alternating optimization, graph theory.
@@ I Introduction
On the other hand, movable antenna (MA) (also known as fluid antenna system (FAS) for antenna position adjustment [9]) has drawn great interest in academia and industry recently. Similar to the IRS, MA can reconfigure wireless channel conditions by enabling multiple antennas to be flexibly moved within a confined region at the transmitter or receiver [10, 11, 12]. The antenna position optimization for MAs has been investigated in various scenarios, such as flexible beamforming [13, 14, 15], physical-layer security [16, 17], multiple-input multiple-output (MIMO) system [18], over-the-air computation [19], cognitive radio [20], integrated sensing and communication (ISAC) [21, 22], etc. Particularly, most of the existing works applied the gradient-based algorithms, e.g., the successive convex approximation (SCA) technique [13, 16, 18, 23] and gradient ascend/descent method [17, 24], to optimize the antenna positions. In addition, some other works proposed a variety of evolutionary algorithms to optimize the antenna positions, such as the particle swarm optimization (PSO) algorithm and its derivative algorithms [19, 25, 26]. Furthermore, the deep learning-based method was introduced in [27] for MA-assisted multicasting, where the feedforward neural networks (FNNs) were utilized to extract useful features from the input direction information and jointly optimize the beamforming and antenna positions. Instead of optimizing the antenna positions in a continuous space, the authors of [28, 20, 29] discretized the movable region into a multitude of sampling points and proposed a graph-based method to select an optimal subset of sampling points for maximizing the communication rates. However, all of the above position optimization algorithms rely on global channel state information (CSI) at any position within the transmit/receive movable region. To facilitate the practical implementation, the authors of [30] proposed a CSI-free MA position optimization approach via the zero-order gradient approximation method.
@@ II System Model and Problem Formulation
Let $\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})\in\mathbb{C}^{M\times N}$ and $\boldsymbol{h}_{\text{IU},k}\in\mathbb{C}^{M\times 1}$ denote the BS-IRS and IRS-user $k$ channels, respectively. Note that the BS-IRS channel depends on the APV, i.e., $\boldsymbol{T}$, while $\boldsymbol{h}_{\text{IU},k}$ is regardless of it. As such, the cascaded BS-IRS-user $k$ channel can be expressed as $\boldsymbol{h}_{k}^{H}(\boldsymbol{T},\boldsymbol{\Phi})=\boldsymbol{h}_{\text{IU},k}^{H}\boldsymbol{\Phi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})\in\mathbb{C}^{1\times N}$, $\forall k\in\mathcal{K}$. Let $\boldsymbol{w}_{k}\in\mathbb{C}^{N\times 1}$ denote the BS’s transmit precoding for user $k$ with $\sum_{k\in\mathcal{K}}{||\boldsymbol{w}_{k}||_{2}^{2}}\leq P$, where $P$ is the BS’s transmit power. Then, the received signal at user $k$ is given by
@@
$R_{k}(\boldsymbol{W},\boldsymbol{\Phi},\boldsymbol{T})=\log_{2}\left(1+\frac{\left|\boldsymbol{h}_{k}^{H}(\boldsymbol{\Phi},\boldsymbol{T})\boldsymbol{w}_{k}\right|^{2}}{\sum_{i\neq k}{\left|\boldsymbol{h}_{k}^{H}(\boldsymbol{\Phi},\boldsymbol{T})\boldsymbol{w}_{i}\right|^{2}}+\sigma^{2}}\right),$ (2)
where $\boldsymbol{W}=[\boldsymbol{w}_{1},\boldsymbol{w}_{2},\cdots,\boldsymbol{w}_{K}]\in\mathbb{C}^{N\times K}$ denotes the transmit precoding matrix. Our objective is to maximize the sum-rate of the $K$ users, by jointly optimizing the BS’s transmit precoding matrix $\boldsymbol{W}$, the IRS’s reflection matrix $\boldsymbol{\Phi}$ and the APV $\boldsymbol{T}$. Hence, the optimization problem is formulated as [^2] Note that the sum-rate maximization problem can be viewed as a dual of the bit-error rate (BER) minimization problem by introducing a BER-related penalty term to the SINRs within the logarithm [40]. Moreover, in the single-user case, the rate maximization becomes equivalent to the BER minimization, since the BER decreases with increasing the user’s achievable rate or SNR.
@@
$||\boldsymbol{t}_{i}-\boldsymbol{t}_{j}||_{2}\geq D_{\min},\forall i,j\in\mathcal{N},i\neq j,$ (3c)
@@
where $D_{\min}$ denotes the minimum spacing between any two MAs to avoid mutual coupling. To characterize the performance limit and ease the performance analysis, we assume that all required CSI is available by applying the existing channel estimation techniques dedicated to MAs/IRSs, e.g., compressed sensing [41, 42]. For example, a viable solution is by equipping the IRS with a few active sensors, such that the BS-IRS channel (for all MA positions) and IRS-user channel can be separately estimated. In addition, we assume that the antenna movement delay is negligible compared to the channel coherence time, which usually holds in various slow-varying scenarios, e.g., smart homes and factories. Notably, the movement delay can be further reduced by employing advanced electronically driven solutions for MAs [10], which also helps prevent potential inter-antenna collisions associated with physical movement.
@@ III-A Performance Analysis
$G_{u}(\boldsymbol{t})=\left(\frac{\lambda}{4\pi}\right)^{2}\left(\left|\sum_{m_{y}=1}^{M_{y}}\sum_{m_{z}=1}^{M_{z}}\frac{|h_{\text{IU},m_{y},m_{z}}|}{D(\boldsymbol{t},m_{y},m_{z})}\right|\right)^{2},$ (9)
@@ III-A1 Optimal Antenna Position
$h_{\text{BU}}(\boldsymbol{t})=\frac{\lambda}{4\pi||\boldsymbol{t}-\boldsymbol{q}_{U}||_{2}}e^{j\frac{2\pi}{\lambda}||\boldsymbol{t}-\boldsymbol{q}_{U}||_{2}},$ (14)
@@
$\boldsymbol{t}^{\star}=\arg\underset{\boldsymbol{t}\in\mathcal{C}_{t}}{\min}\,\,||\boldsymbol{t}-\boldsymbol{q}_{U}||_{2}.$ (15)
@@ III-A2 Fluctuation of Channel Power Gain
On the other hand, we analyze the effects of the BS-IRS distance on (17). For simplicity, we assume that the MA array is perpendicular to the IRS. Then, $R(\boldsymbol{t}_{i})$ can be expressed as $R(\boldsymbol{t}_{i})=\sqrt{d_{\text{BI}}^{2}+d^{2}(\boldsymbol{t}_{i})}$, $i=1,2$, where $d(\boldsymbol{t}_{i})=||\boldsymbol{q}_{B}-\boldsymbol{t}_{i}||_{2}$ denotes the distance between the $i$-th location and the center of the MA array, and $d_{\text{BI}}$ denotes the distance from the center of $\mathcal{C}_{t}$ to that of the IRS. As $R(\boldsymbol{t}_{1})<R(\boldsymbol{t}_{2})$, $H(M_{y},M_{z},\boldsymbol{t}_{1})\pm H(M_{y},M_{z},\boldsymbol{t}_{2})$ monotonically decreases with $d_{\text{BI}}$. Thus, $G_{u}(\boldsymbol{t}_{1})-G_{u}(\boldsymbol{t}_{2})$ also decreases with $d_{\text{BI}}$. This indicates that a more/less significant fluctuation of (10) will be resulted if the BS-IRS distance reduces/increases.
@@ III-A3 Far-Field BS-IRS Channel
All of the above analytical results are derived under the general NUSW-based BS-IRS channel model. Next, we consider a special case with the far-field BS-IRS channel. In this case, the movement region is approximated as a point source; thus, the distance between the MA and the $(m_{y},m_{z})$-th IRS reflecting element in (18) can be approximated as identical, i.e.,
@@
$\boldsymbol{w}(\boldsymbol{T})=\frac{\boldsymbol{v}(\boldsymbol{T})}{||\boldsymbol{v}(\boldsymbol{T})||_{2}},$ (19)
@@ III-B Proposed Solution to (P2)
$\boldsymbol{w}(\boldsymbol{T},\boldsymbol{\Phi})=\frac{\left(\boldsymbol{h}_{\text{IU}}^{H}\boldsymbol{\Phi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})\right)^{H}}{||\boldsymbol{h}_{\text{IU}}^{H}\boldsymbol{\Phi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})||_{2}}.$ (20)
@@
(P2) $\underset{\boldsymbol{\Phi},\boldsymbol{T}}{\max}\quad\gamma(\boldsymbol{T},\boldsymbol{\Phi})=||\boldsymbol{h}_{\text{IU}}^{H}\boldsymbol{\Phi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})||^{2}_{2}$
@@ III-B1 Optimizing $\boldsymbol{\Phi}$ for a Given $\boldsymbol{T}$
(P2-1) $\underset{\boldsymbol{\varphi}}{\max}\quad\gamma(\boldsymbol{T},\boldsymbol{\Phi})=\left|\left|\sum_{m=1}^{M}{\boldsymbol{g}_{1,m}e^{j\varphi_{m}}}\right|\right|^{2}_{2}$
@@
which, however, is still a non-convex optimization problem. To tackle this non-convexity, we first define $\boldsymbol{\alpha}_{m}=\sum_{i\neq m}^{M}{\boldsymbol{g}_{1,i}e^{j\varphi_{i}}}$, $\forall m\in\mathcal{M}$, and rewrite the objective function in (P2-1) as a more tractable form as
$\gamma(\boldsymbol{T},\boldsymbol{\Phi})$ $=\left|\left|\boldsymbol{g}_{1,m}e^{j\varphi_{m}}+\boldsymbol{\alpha}_{m}\right|\right|^{2}_{2}$ (23)
$=\left(\boldsymbol{g}_{1,m}e^{j\varphi_{m}}+\boldsymbol{\alpha}_{m}\right)\left(\boldsymbol{g}_{1,m}e^{j\varphi_{m}}+\boldsymbol{\alpha}_{m}\right)^{H}$
$=\left|\left|\boldsymbol{g}_{1,m}\right|\right|^{2}_{2}+\left|\left|\boldsymbol{\alpha}_{m}\right|\right|^{2}_{2}+2\Re\left\{\boldsymbol{\alpha}_{m}\boldsymbol{g}_{1,m}^{H}e^{-j\varphi_{m}}\right\}.$
Next, we propose to utilize an element-wise BCD method to optimize the IRS reflecting coefficients sequentially. Specifically, it can be easily shown from (23) that for any given $\varphi_{i}$, $i\in{\cal M}$, $i\neq m$, the optimal $\varphi_{m}$ for (P2-1) should be such that the phase of $\boldsymbol{\alpha}_{m}\boldsymbol{g}_{1,m}^{H}$ and $e^{-j\varphi_{m}}$ are identical, which is given by
$\varphi_{m}^{\star}=\arg\left(\boldsymbol{\alpha}_{m}\boldsymbol{g}_{1,m}^{H}\right).$ (24)
As such, in the $m$-th BCD iteration, we can fix $\varphi_{i}$, $i\neq m$, $i\in{\cal M}$ and optimize $\varphi_{m}$ as (24). Then, the BCD iteration proceeds until the phase shifts of all $M$ IRS reflecting elements have been updated.
@@ IV-B1 Optimizing $\boldsymbol{W}$ for Given $\mathbf{\Phi}$ and $\boldsymbol{T}$
where $\mu\geq 0$ is the optimal dual variable that ensures $\sum_{k=1}^{K}\left\|\boldsymbol{w}_{k}\right\|_{2}^{2}\leq P$. Note that $\sum_{k=1}^{K}\left\|\boldsymbol{w}_{k}\right\|_{2}^{2}$ is a monotonically decreasing function w.r.t. $\mu$ [34]. Thus, we can find $\mu$ via a bisection search.
@@ V Numerical Results
    Antenna selection (AS): In this benchmark, $A/D_{\min}$ FPAs are deployed within $\mathcal{C}_{t}$ and separated by the minimum distance $D_{\min}$. Among them, $N$ antennas are selected for transmission. The associated optimization problem can be solved by applying a similar AO algorithm as in Section IV-B by setting $\delta_{s}=D_{\min}$.
  3. 3.
@@ V-A Single-User System
First, we consider the single-user setup and plot in Fig. 3 the achievable rate at the user versus the BS-IRS distance, with $N=4$ and $L=8$. The IRS-user distance is fixed as $d_{\text{IU}}=30$ m. It is observed that our proposed algorithm achieves higher received SNRs than all other benchmarks, thanks to channel reconfiguration abilities for both MAs and IRSs. However, the performance gain of our proposed algorithm over the FPA benchmark is marginal and decreases with the BS-IRS distance, which shows a similar trend to the case with the LoS BS-IRS channel in Fig. 2(b). In contrast, if the IRS phase shifts are configured randomly, employing MAs can yield more significant performance gain over the conventional FPAs. For example, when the BS-IRS distance is 3 m, the performance gain of the proposed algorithm over the FPA benchmark is only 0.27 bps/Hz (or 2.53%); while that of the MA-RPS benchmark over the FPA-RPS benchmark is 1.30 bps/Hz (or 36.94%). This is mainly attributed to the channel reconfiguration from the IRS. More specifically, under the optimized IRS reflection, the correlations among all transmit paths between the BS and the IRS may be decreased, which also reduces the spatial diversity gain reaped from the multi-path effect. In addition, the lack of performance gains over the FPA benchmark may also be attributed to the structural similarity between the IRS and field-response channels, as observed from (1) and (52), where the path gain of each path plays a role similar to the reflection coefficient of each IRS element. This endows the field-response channel with reconfigurability, thus diminishing the performance gain by the MAs. Due to the above reasons, the performance gain of MAs over the AS benchmark is observed to be even more marginal compared to the FPA benchmark. In contrast, under the random IRS reflection, the IRS can be simply treated as a scatterer in the environment, thus barely compromising the multi-path effect.
To validate the above claim, we plot in Figs. 4(a) and 4(b) the channel power gains from the BS to the user within $\mathcal{C}_{t}$ by the proposed algorithm and the MA-RPS benchmark, respectively. It is observed that compared to the proposed algorithm, the MA-RPS benchmark yields more local maxima and minima, i.e., more significant fluctuation, within the transmit array. Furthermore, the gap between the maximum and the minimum channel power gains in the MA-RPS benchmark is around 15.62 dB, which far exceeds that in the proposed algorithm, i.e., 3.48 dB. This may be attributed to the fact that the movement region of MAs is typically limited to a wavelength level. As a result, the IRS’s passive beamforming may suffice to ensure the channel power gain across all MA positions within the movement region. This effect is particularly evident in the far-field scenario, where the point-source approximation is applied to the movement region, as discussed in Section III-A. This observation validates our previous claim regarding the effects of the random versus optimized IRS reflections.